​​​​​​1. An important part of the customer service responsibilities of a cable company relates to the speed with which trouble in service can be repaired. Historically, the data show that the likelihood is 0.85 that troubles in a residential service can be repaired on the same day. Use the binomial formula to calculate.

c. For the first five troubles reported on a given day, what is the probability that: At least 3 troubles will be repaired on the same day.
Answer #1

Binomial Distribution

If 'X' is the random variable representing the number of successes, the probability of getting ‘r’ successes and ‘n-r’ failures, in 'n' trails, ‘p’ probability of success ‘q’=(1-p) is given by the probability function

P q

p : Probability that a trouble will be repaired on the same day = 0.85

q = 1-p = 1-0.85=0.15

n : Number of troubles reported on a given day : n= 5

X : Number of troubles will be repaired on the same day.

X follows a Binomial with n= 5 and p =0.85

And the probability function of X is

P(X = r) = ( ) (0.85) (0.15)5-r

Probability that: At least 3 troubles will be repaired on the same day = P(X\small \geq3) = P(X=3)+P(X=4)+P(X=5)

P(X = 3) = ( 2) × 0.853 x 0.15 x 0.85 x 0.152-10 x 0.614125 x 0.02250.1382 10 x 0.614125 x 0.0225 0.1382

PlX = 4) = (·) x 0.854 x 0.15-5 × 0.52200625 x 0.15 0.3915

x 0.855 x 0.150-1 x 0.4437053 125 x 1 0.4437

P(X\small \geq3) = P(X=3)+P(X=4)+P(X=5) = 0.1382 + 0.3915 + 0.4437 = 0.9734

Probability that: At least 3 troubles will be repaired on the same day = 0.9734